A SUV can be bought for $28,000, or for $2500 down and $10000 at the end of each year for three years. If the money can earn 7% compounded quarterly, determine which deal is better.
(first option; present value of the second option is $28,654.28)
Answer #1

Present value of 2nd option = cash down payment + PV of three installments

Effective interest rate = (1 + 7%/4)^4 - 1 = 7.186%

= 2500 + 10000 / (1 + 7.186%)^1 + 10000 / (1 + 7.186%)^2 + 10000 / (1+7.186%)^3

= 28654.23

Since we have to pay 28000 today in the first option, it is better to avail it as the present value of Option 2 is higher